On Jeff's outgoing trip, he travels 648 miles and uses 18 gallons of gas. On his return trip, Jeff travels a different route. Jeff travels 672 miles on the return trip and uses 24 gallons of gas.
How many miles per gallon did Jeff's car get on the entire trip?
23 miles per gallon
28 miles per gallon
36 miles per gallon
32.5 miles per gallon
31.4 miles per gallon

Answer by Theo(13342)

going he used 18 gallons for 648 miles.
coming back he used 24 gallons for 672 miles.
on the trip going and coming back he used 18 + 24 = 42 gallons for 648 + 672 = 1320 miles.
miles divided by gallons = miles per gallon, therefore 1320/42 = 31.42857143 miles per gallon.
that's for the whole trip.
that should be your answer.